# Interference Among First-Price Pacing Equilibria: A Bias and Variance Analysis

### v1

@@ 4.3. Asymptotic Normality and Inference
4.4. Results for Revenue
The revenue in the good market is defined as ${\mathsf{{REV}}}^{*}=\int\max_{i}\{{v_{i}(\theta)}{\beta^{*}_{i}}\}s(\theta)\mathop{}\!\mathrm{d}\theta$. We define the debiased revenue surrogate and its estimator
(9) $\widetilde{{\mathsf{{REV}}}}^{*}=\int\max_{i}\{{v_{i}(\theta)}\widetilde{\beta}^{*}_{i}\}s(\theta)\mathop{}\!\mathrm{d}\theta\;,$
(10) $\widehat{{\mathsf{{REV}}}}=\frac{1}{1-\alpha}\frac{1}{t}{\sum_{\tau=1}^{t}}1_{\Theta_{\mathsf{good}}}({\theta^{\tau}})\max_{i}\{{v_{i}(\theta^{\tau})}\widehat{\beta}_{i}\}\;.$
In words, we approximate ${\mathsf{{REV}}}^{*}$ with the revenue generated in the market where buyers bid according to the pacing profile $\widetilde{\beta}^{*}$, and then we estimate this via plug-in estimation on the subset of items that are good.
Theorem 4 ().
@@ 4.4. Results for Revenue
We have two methods to construct CIs for the debiased revenue surrogate in Eq. 7. The first method is based on the asymptotic normality results in Thm. 4. In the second method, a revenue CI is constructed based on a CI for $\beta^{*}_{\alpha}$. Suppose ${\small\mathbb{{CI}}}$ is some confidence region for $\beta^{*}_{\alpha}$. Then $\{\frac{1}{1-\alpha}\frac{1}{t}{\sum_{\tau=1}^{t}}\allowbreak 1_{\Theta_{\mathsf{good}}}({\theta^{\tau}})\max_{i}{v_{i}(\theta^{\tau})}{\beta_{i}}:\beta\in{\small\mathbb{{CI}}}\}$ serves as a CI for revenue. Another way to utilize a pacing multiplier CI is based on the observation that revenue is a monotone function of the pacing multiplier. Let $\underline{{\beta_{i}}}=\min\{{\beta_{i}}:\beta\in{\small\mathbb{{CI}}}\}$ and $\bar{\beta}_{i}=\max\{{\beta_{i}}:\beta\in{\small\mathbb{{CI}}}\}$. Then another natural (potentially wider) CI for $\widetilde{{\mathsf{{REV}}}}^{*}$ is $[\underline{{\mathsf{{REV}}}},\overline{{\mathsf{{REV}}}}]$, where
(11) $\underline{{\mathsf{{REV}}}}=\frac{1}{1-\alpha}\frac{1}{t}{\sum_{\tau=1}^{t}}1_{\Theta_{\mathsf{good}}}({\theta^{\tau}})\max_{i}{v_{i}(\theta^{\tau})}\underline{{\beta_{i}}}\;,\quad\overline{{\mathsf{{REV}}}}=\frac{1}{1-\alpha}\frac{1}{t}{\sum_{\tau=1}^{t}}1_{\Theta_{\mathsf{good}}}({\theta^{\tau}})\max_{i}{v_{i}(\theta^{\tau})}\bar{\beta}_{i}\;.$
5. Parallel A/B Test under Interference
@@
Step 1. The treatment effects. Let ${\mathscr{M}}(w,k)={\small\mathsf{{FPPE}}}(\{b_{i}\}_{i\in C_{k}},\{v_{i}({\cdot},w)\}_{i\in C_{k}},s_{k},\Theta_{k})$be the $k$-th limit submarket under treatment $w\in\{0,1\}$, and let revenue be ${\mathsf{{REV}}}^{*}_{k}(w)$. Then the treatment effect in the submarket $k$ is $\tau^{*}_{k}={\mathsf{{REV}}}^{*}_{k}(1)-{\mathsf{{REV}}}^{*}_{k}(0)$. Equivalently, all $K$ submarkets can be formed simultaneously, ${\mathscr{M}}(w)={\small\mathsf{{FPPE}}}\left(\frac{1}{K}b,v(w),s,\Theta\right).$ The scaling $1/K$ in budgets ensures that a buyer in $C_{k}$ has budget $b_{i}/K$ to bid for items in $\Theta_{k}$, whose supply in ${\mathscr{M}}(w)$ is $\frac{1}{K}s_{k}$. The $k$-th component of ${\mathscr{M}}(w)$ corresponds to ${\mathscr{M}}(w,k)$.
@@
The data observed following the described A/B test scheme can be compactly represented by two finite FPPEs (one for each treatment) (Def. 4): ^$\widehat{\mathscr{M}}_{\alpha}(w)={\small\widehat{\textsf{{FPPE}}}}(\pi_{w}tb,v(w),1,\gamma)$ , where $\gamma\subset\Theta$ consists of $t_{w}=\lfloor\pi_{w}tK\rfloor$ i.i.d. draws from the mixture $\alpha s^{\prime}+(1-\alpha)s$.
It can be shown that the observed market ^$\widehat{\mathscr{M}}_{\alpha}(w)$ converges to ${\mathscr{M}}_{\alpha}(w)={\small\mathsf{{FPPE}}}\ \left(\frac{1}{K}b,v(w),s_{\alpha},\Theta\right).$
Step 3. Debiasing $\beta$. Following previous section, we can debias ^$\widehat{\mathscr{M}}_{\alpha}(w)$ to approximate ${\mathscr{M}}(w)$. Let $\widehat{\beta}(w)$ be the debiased pacing multiplier.
Step 4. Revenue estimator. Estimate $\widehat{\mathsf{{REV}}}_{k}(w)$ by
@@
Then $\widehat{\tau}_{k}=\widehat{\mathsf{{REV}}}_{k}(1)-\widehat{\mathsf{{REV}}}_{k}(0)$ and a variance estimate can be obtained by either asymptotic normality, confidence region of $\beta$, or bootstrap.
6. Experiment
To evaluate our proposed framework and debiased estimator, we run both fully synthetic and semi-synthetic simulations to check if the proposed estimator for beta and revenue are indeed less biased as well as the coverage of the proposed estimator.
6.1. Fully synthetic experiments
@@
In the fully synthetic experiments, we simulate 10 buyers and 1000 good items in two submarkets, with a varying number of bad items (up to 500) in order to study the effect of the contamination parameter $\alpha$. For each $\alpha$, we randomly sample a budget from a uniform distribution for each buyer, and compute $\beta^{*}$ and ${\mathsf{{REV}}}^{*}$ from the limit pure market ${\mathscr{M}}_{0}$ with a uniformly distributed value function in each submarket. To check if the debiased surrogate reduces bias, we compute $\beta^{*}_{\alpha}$ and ${\mathsf{{REV}}}^{*}_{\alpha}$ from the limit market with interference ${\mathscr{M}}_{\alpha}$ and their surrogates, $\widetilde{\beta}^{*}$ and $\widetilde{\mathsf{{REV}}}^{*}$. We look at the normalized bias for the surrogate, defined as $\|\widetilde{\beta}^{*}-\beta^{*}\|_{2}/\|\beta^{*}\|_{2}$ for pacing multipliers and as $|\widetilde{\mathsf{{REV}}}^{*}/{\mathsf{{REV}}}^{*}-1|$ for revenue, and similarly defined for the limit quantities. Fig 4 shows the normalized bias curves as a function of $\alpha$. It can be seen that the magnitude of the bias increases with $\alpha$, for both the variables in the limit market with interference ${\mathscr{M}}_{\alpha}$ and their debiased surrogates. We see that the bias of the debiased surrogates is indeed much smaller than the contaminated limit quantities.
@@
Figure 4. Normalized bias (in percent of true value) as a function of $\alpha$ in fully synthetic experiments. $\widetilde{\beta}^{*}$ and $\widetilde{\mathsf{{REV}}}^{*}$ are the debiased surrogates for pacing multiplier and revenue in the limit market with interference ${\mathscr{M}}_{\alpha}$.
@@
Next, we check the coverage of the proposed variance estimator. For each $\alpha$ and each budget sample, we run 100 simulations in the following way: We sample items (or their values for each buyer) considering two submarkets and bad items. We then run the finite FPPE with bad items and obtain a baseline estimate for pacing multiplier and revenue without applying the debiasing procedure. Then, we apply the debiasing procedure to compute the debiased estimates. For each simulation, we check if the debiased surrogate is within the confidence interval of the debiased estimator. Finally, we aggregate them to compute the estimated coverage of the estimator. The results for both pacing multiplier and revenue are shown in Table 3. For the coverage of $\widehat{\beta}$, we first compute the coverage of each component and report only the average in the table. It is shown that $\widehat{\beta}$ is slightly over-covered. We found that the over-coverage is mainly due to the over-estimation of the estimator’s variance. For revenue, we construct the CI using the two approaches as mentioned in Sec. 4.4: one based on Eq. 11 and the other using parametric bootstrap based on the estimated asymptotic distribution of $\widehat{\beta}$ (with ”(b)” in the column names). First of all, we show in Fig 5 that both CIs converges to the true value from the limit market with interfrence ${\mathscr{M}}_{\alpha}$ as the number of items goes to infinity. Then, in  Table 3, we show that the bootstrap CI width is much smaller compared to that from Eq. 11. However, both CIs are over-covered, higher than the nominal level (95%). It is expected though that the CI from Eq. 11 is relatively large and hence conservative since it uses the min and max of each $\beta_{i}$, hence likely living outside of the confidence region ($n$-dimension ellipsoid) for $\widehat{\beta}$ and leading to over-coverage. For the bootstrap CI, it requires enough bootstrap samples to fully explore the whole boundary of the confidence region of $\widehat{\beta}$, as well as an accurate estimate of the asymptotic distribution of $\widehat{\beta}$. In our fully synthetic simulation, our hypothesis is that the over-coverage of the bootstrap CI is mainly due to the over-estimation of the variance for $\widehat{\beta}$.
$\widehat{\beta}$ $\widehat{\mathsf{{REV}}}$
@@
Table 3. Coverage of $\widehat{\beta}$ and revenue estimates in fully synthetic experiments. All quantities are averaged over 100 simulations for each $\alpha$. The coverage of $\widehat{\beta}$ is averaged over all components of $\widehat{\beta}$. For revenue estimates, the columns with ”(b)” represent the quantities from the bootstrap CI, while the columns without ”(b)” are for the CI from Eq. 11. The CI widths are normalized by the revenue from the limit market ${\mathscr{M}}_{\alpha}$. $\alpha$ is expressed as the ratio of the number of bad items and the total items.

### v3

@@ Appendix E Results for Revenue
The revenue in the good market is defined as ${{{REV}}}^{*}=\int\max_{i}\{{v_{i}(\theta)}{\beta^{*}_{i}}\}s(\theta)\mathop{}\!\mathrm{d}\theta$. We define the debiased revenue surrogate and its estimator
$\widetilde{{{{REV}}}}^{*}=\int\max_{i}\{{v_{i}(\theta)}\widetilde{{\bm{\beta}}}^{*}_{i}\}s(\theta)\mathop{}\!\mathrm{d}\theta\;,$ (20)
$\widehat{{{{REV}}}}=\frac{1}{1-\alpha}\frac{1}{t}{\sum_{\tau=1}^{t}}1_{\Theta_{\mathsf{good}}}({\theta^{\tau}})\max_{i}\{{v_{i}(\theta^{\tau})}\widehat{\bm{\beta}}_{i}\}\;.$ (21)
In words, we approximate ${{{REV}}}^{*}$ with the revenue generated in the market where buyers bid according to the pacing profile $\widetilde{{\bm{\beta}}}^{*}$, and then we estimate this via plug-in estimation on the subset of items that are good.
Theorem 4.
@@
We have two methods to construct CIs for the debiased revenue surrogate in Eq. 7. The first method is based on the asymptotic normality results in Thm. 4. In the second method, a revenue CI is constructed based on a CI for ${\bm{\beta}}^{*}_{\alpha}$. Suppose ${\mathbb{{CI}}}$ is some confidence region for ${\bm{\beta}}^{*}_{\alpha}$. Then $\{\frac{1}{1-\alpha}\frac{1}{t}{\sum_{\tau=1}^{t}}\allowbreak 1_{\Theta_{\mathsf{good}}}({\theta^{\tau}})\max_{i}{v_{i}(\theta^{\tau})}{\beta_{i}}:{\bm{\beta}}\in{\mathbb{{CI}}}\}$ serves as a CI for revenue. Another way to utilize a pacing multiplier CI is based on the observation that revenue is a monotone function of the pacing multiplier. Let $\underline{{\beta_{i}}}=\min\{{\beta_{i}}:{\bm{\beta}}\in{\mathbb{{CI}}}\}$ and $\bar{{\bm{\beta}}}_{i}=\max\{{\beta_{i}}:{\bm{\beta}}\in{\mathbb{{CI}}}\}$. Then another natural (potentially wider) CI for $\widetilde{{{{REV}}}}^{*}$ is $[\underline{{{{REV}}}},\overline{{{{REV}}}}]$, where
$\underline{{{{REV}}}}=\frac{1}{t}{\sum_{\tau=1}^{t}}\frac{1_{\Theta_{\mathsf{good}}}({\theta^{\tau}})}{1-\alpha}\max_{i}{v_{i}(\theta^{\tau})}\underline{{\beta_{i}}}\;,\quad\overline{{{{REV}}}}=\frac{1}{t}{\sum_{\tau=1}^{t}}\frac{1_{\Theta_{\mathsf{good}}}({\theta^{\tau}})}{1-\alpha}\max_{i}{v_{i}(\theta^{\tau})}\bar{{\bm{\beta}}}_{i}\;.$ (22)
Appendix F Parallel A/B Test under Interference
@@
Step 1. The treatment effects. Let ${\mathscr{M}}(w,k)={\mathsf{{FPPE}}}(\{b_{i}\}_{i\in C_{k}},\{v_{i}({\cdot},w)\}_{i\in C_{k}},s_{k},\Theta_{k})$be the $k$-th limit submarket under treatment $w\in\{0,1\}$, and let revenue be ${{{REV}}}^{*}_{k}(w)$. Then the treatment effect in the submarket $k$ is $\tau^{*}_{k}={{{REV}}}^{*}_{k}(1)-{{{REV}}}^{*}_{k}(0)$. Equivalently, all $K$ submarkets can be formed simultaneously, ${\mathscr{M}}(w)={\mathsf{{FPPE}}}\left(\frac{1}{K}\bm{b},\bm{v}(w),s,\Theta\right).$ The scaling $1/K$ in budgets ensures that a buyer in $C_{k}$ has budget $b_{i}/K$ to bid for items in $\Theta_{k}$, whose supply in ${\mathscr{M}}(w)$ is $\frac{1}{K}s_{k}$. The $k$-th component of ${\mathscr{M}}(w)$ corresponds to ${\mathscr{M}}(w,k)$.
@@
The data observed following the described A/B test scheme can be compactly represented by two finite FPPEs (one for each treatment) (Def. 5): ^$\widehat{\mathscr{M}}_{\alpha}(w)={\widehat{\textsf{{FPPE}}}}(\pi_{w}t\bm{b},\bm{v}(w),1,\gamma)$ , where $\gamma\subset\Theta$ consists of $t_{w}=\lfloor\pi_{w}tK\rfloor$ i.i.d. draws from the mixture $\alpha s^{\prime}+(1-\alpha)s$.
It can be shown that the observed market ^$\widehat{\mathscr{M}}_{\alpha}(w)$ converges to ${\mathscr{M}}_{\alpha}(w)={\mathsf{{FPPE}}}\ \left(\frac{1}{K}\bm{b},\bm{v}(w),s_{\alpha},\Theta\right).$
Step 3. Debiasing ${\bm{\beta}}$. Following previous section, we can debias ^$\widehat{\mathscr{M}}_{\alpha}(w)$ to approximate ${\mathscr{M}}(w)$. Let $\widehat{\bm{\beta}}(w)$ be the debiased pacing multiplier.
Step 4. Revenue estimator. Estimate $\widehat{{{REV}}}_{k}(w)$ by
@@
Then $\widehat{\tau}_{k}=\widehat{{{REV}}}_{k}(1)-\widehat{{{REV}}}_{k}(0)$ and a variance estimate can be obtained by either asymptotic normality, confidence region of ${\bm{\beta}}$, or bootstrap.
Appendix G Heuristic Derivation of the Debiased Surrogate
We present a heuristic argument for Eq. 6 here; for a rigorous treatment we need a perturbation theory result which is given in lemma 1 in the appendix. Define the contaminated EG objective $H_{\alpha}({\bm{\bm{\beta}}})=\int F({\bm{\bm{\beta}}},\theta)s_{\alpha}\mathop{}\!\mathrm{d}\theta$. Then ${\bm{H}}_{\alpha}=\nabla^{2}H_{\alpha}({\bm{\beta}}^{*}_{\alpha})$. The map $({\bm{\bm{\beta}}},\alpha)\mapsto H_{\alpha}({\bm{\bm{\beta}}})$ has gradient $[\nabla H({\bm{\beta}}^{*}),\int F({\bm{\beta}}^{*}_{\alpha},\theta)(s^{\prime}-s)\mathop{}\!\mathrm{d}\theta]^{\scriptscriptstyle\mathsf{T}}$ and Hessian $[{\bm{H}}_{\alpha},-{\bm{\delta}}_{\alpha};-{\bm{\delta}}_{\alpha}^{\scriptscriptstyle\mathsf{T}},0]$ at $({\bm{\beta}}^{*}_{\alpha},\alpha)$. Then we have the following quadratic approximation of $H_{\alpha-\epsilon}({\bm{\bm{\beta}}})$
$H_{\alpha-\epsilon}({\bm{\bm{\beta}}})\approx H_{\alpha}({\bm{\beta}}^{*}_{\alpha})+\begin{pmatrix}{\bm{\bm{\beta}}}-{\bm{\beta}}^{*}_{\alpha}\\ -\epsilon\end{pmatrix}^{\scriptscriptstyle\mathsf{T}}\begin{pmatrix}\nabla H_{\alpha}({\bm{\beta}}^{*}_{\alpha})\\ \int F({\bm{\beta}}^{*}_{\alpha},\theta)(s^{\prime}-s)\mathop{}\!\mathrm{d}\theta\end{pmatrix}$
$+\frac{1}{2}\begin{pmatrix}{\bm{\bm{\beta}}}-{\bm{\beta}}^{*}_{\alpha}\\ -\epsilon\end{pmatrix}^{\scriptscriptstyle\mathsf{T}}\begin{pmatrix}{\bm{H}}_{\alpha}&-{\bm{\delta}}_{\alpha}\\ -{\bm{\delta}}_{\alpha}^{\scriptscriptstyle\mathsf{T}}&0\end{pmatrix}\begin{pmatrix}{\bm{\bm{\beta}}}-{\bm{\beta}}^{*}_{\alpha}\\ -\epsilon\end{pmatrix}\;.$
And so
${\bm{\beta}}^{*}_{\alpha-\epsilon}=\operatorname*{arg\,min}_{{\bm{\beta}}\in(0,1]^{n}}H_{\alpha-\epsilon}({\bm{\bm{\beta}}})\approx\operatorname*{arg\,min}_{{\bm{\beta}}\in(0,1]^{n}}(\nabla H_{\alpha}({\bm{\beta}}^{*}_{\alpha})+\epsilon{\bm{\delta}}_{\alpha})^{\scriptscriptstyle\mathsf{T}}{\bm{\bm{\beta}}}$
$+\tfrac{1}{2}({\bm{\bm{\beta}}}-{\bm{\beta}}^{*}_{\alpha})^{\scriptscriptstyle\mathsf{T}}{\bm{H}}_{\alpha}({\bm{\bm{\beta}}}-{\bm{\beta}}^{*}_{\alpha})\;,$
where we dropped terms that do not involve ${\bm{\bm{\beta}}}$.
Next we need two observations. For $\epsilon$ small enough, we have ${\bm{\beta}}^{*}_{\alpha-\epsilon,i}=1$ if $\beta^{*}_{\alpha,i}=1$. Intuitively, buyers in the market ${\mathscr{M}}_{\alpha}$ with $\beta^{*}_{\alpha,i}=1$ have some nonzero leftover budget, which acts as a buffer to small changes in supply, and thus their pacing multipliers remain one. Next, we also note ${\bm{\beta}}\mapsto\nabla H_{\alpha}({\bm{\beta}}^{*}_{\alpha})^{\scriptscriptstyle\mathsf{T}}{\bm{\bm{\beta}}}$ is constant on the set $\{{\bm{\bm{\beta}}}:{\beta_{i}}=1\text{ if }\beta^{*}_{\alpha,i}=1\}$ by complementary slackness. Then we further simplify and obtain
${\bm{\beta}}^{*}_{\alpha-\epsilon}\approx\operatorname*{arg\,min}_{{\beta_{i}}=1\text{ if }\beta^{*}_{\alpha,i}=1}\epsilon{\bm{\delta}}_{\alpha}^{\scriptscriptstyle\mathsf{T}}{\bm{\bm{\beta}}}+\tfrac{1}{2}({\bm{\bm{\beta}}}-{\bm{\beta}}^{*}_{\alpha})^{\scriptscriptstyle\mathsf{T}}{\bm{H}}_{\alpha}({\bm{\bm{\beta}}}-{\bm{\beta}}^{*}_{\alpha})\;.$
The right-hand side is just a linearly constrained quadratic optimization problem, which admits a closed-form solution ${\bm{\beta}}^{*}_{\alpha}-\epsilon({\bm{\bm{P}}}_{\alpha}{\bm{H}}_{\alpha}{\bm{\bm{P}}}_{\alpha})^{\dagger}{\bm{\delta}}_{\alpha}$.
Appendix H Fully synthetic experiments
@@
In the fully synthetic experiments, we simulate 10 buyers and 1000 good items in two submarkets, with a varying number of bad items (up to 500) in order to study the effect of the contamination parameter $\alpha$. For each $\alpha$, we randomly sample a budget from a uniform distribution for each buyer, and compute ${\bm{\beta}}^{*}$ and ${{{REV}}}^{*}$ from the limit pure market ${\mathscr{M}}_{0}$ with a uniformly distributed value function in each submarket. Fig 5 shows the normalized bias curves as a function of $\alpha$ for the fully synthetic experiment, which is similar to that from semi-synthetic experiments in Fig 3.
@@
Figure 5: Normalized bias (in percent of true value) as a function of $\alpha$ in fully synthetic experiments. $\widetilde{\bm{\beta}}^{*}$ and $\widetilde{{{REV}}}^{*}$ are the debiased surrogates for pacing multiplier and revenue in the limit market with interference ${\mathscr{M}}_{\alpha}$.
@@
We also check the coverage of the proposed variance estimator in fully synthetic experiments. The coverage of $\widehat{\bm{\beta}}$ and revenue are estimated in the same way as in the semi-synthetic ones. It is shown in Fig 6 that both CIs converges to the true value from the limit market with interfrence ${\mathscr{M}}_{\alpha}$ as the number of items goes to infinity. Then, in  Table 2, we show that the bootstrap CI width is much smaller compared to that from Eq. 22. However, unlike in the semi-synthetic experiments, both CIs are over-covered, higher than the nominal level (95%). It is expected though that the CI from Eq. 22 is relatively large and hence conservative since it uses the min and max of each $\beta_{i}$, hence likely living outside of the confidence region ($n$-dimension ellipsoid) for $\widehat{\bm{\beta}}$ and leading to over-coverage. For the bootstrap CI, it requires enough bootstrap samples to fully explore the whole boundary of the confidence region of $\widehat{\bm{\beta}}$, as well as an accurate estimate of the asymptotic distribution of $\widehat{\bm{\beta}}$. In our fully synthetic simulation, our hypothesis is that the over-coverage of the bootstrap CI is mainly due to the over-estimation of the variance for $\widehat{\bm{\beta}}$.
$\widehat{\bm{\beta}}$ $\widehat{{{REV}}}$
@@
Table 2: Coverage of $\widehat{\bm{\beta}}$ and revenue estimates in fully synthetic experiments. All quantities are averaged over 100 simulations for each $\alpha$. The coverage of $\widehat{\bm{\beta}}$ is averaged over all components of $\widehat{\bm{\beta}}$. For revenue estimates, the columns with "(b)" represent the quantities from the bootstrap CI, while the columns without "(b)" are for the CI from Eq. 22. The CI widths are normalized by the revenue from the limit market ${\mathscr{M}}_{\alpha}$. $\alpha$ is expressed as the ratio of the number of bad items and the total items.